OpenStudy (anonymous):

Bayside Insurance offers two health plans. Under A, Giselle would have to pay the first $190 of her medical bills, plus 35% of the rest. Under plan B, Giselle would pay the first $240, but only 20% of the rest. For what amount of medical bills will plan B save Giselle money? Assume she has over $20 in bills. She would save with plan B if she has more than___in bills???

OpenStudy (kropot72):

The break even point for the two plans can be found be solving the following equation for A where A is the amount of a bill: 190 + 0.35(A - 190) = 240 + 0.2(A - 240) Do you follow? If so can you solve for A?

OpenStudy (kropot72):

Collecting terms in A on the left hand side and constant terms on the right hand side gives: 0.15A = 240 - 123.5 - 48 = 68.5 A = $456.67 So Giselle would save with plan B for bills more than $456.67

OpenStudy (anonymous):

Thank you

OpenStudy (kropot72):

You're welcome :)